OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is even and 1 die shows a 3?

OpenStudy (cwrw238):

probability that one die shows a 3 = (1/6) now for the sum to be even the other die must show one of which values

OpenStudy (anonymous):

hmm... 1, 3, 5?

OpenStudy (cwrw238):

exactly so whats probability of 1,3 or 5?

OpenStudy (anonymous):

1/2

OpenStudy (cwrw238):

right now the two events ( the 3 and 1,3,5) are independent so the 2 probabilities are multiplied 1/2 * 1/6

OpenStudy (anonymous):

1/12

OpenStudy (anonymous):

is it 1/12 then?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

wow it's wrong... it's 5/36

OpenStudy (cwrw238):

hmm - lets look at it again

OpenStudy (cwrw238):

i've missed something here - yes the 3 can be on either dice - well if the 3 is on the other dice instead that doubles the probability - giving 1/6 - thats still wrong

OpenStudy (cwrw238):

- oh 2 '3' isn't two events - just one - so that subtracts 1/36 from 1/6 giving 5/36 -this was more complicated than it appears sorry guy - but we got there in the end

OpenStudy (cwrw238):

do you follow? all possibilities are 3 and 1 3 and 5 1 and 3 5 and 3 3 and 3 probabilities for each is 1/36 total probabilities = 5/36 - that way was a lot easier!!!

OpenStudy (cwrw238):

lol